# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v3

@@ 3. Classification of orbits
The sum of the coefficients is exactly 883, the number of supports in the list plus three supports with restrictions.
Since each form from $S$ defines its own orbit, the set $S$ defines $|S|$ orbits. But in work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is calculated using group-theoretic methods. And it is exactly equal to $|S|$. Hence, over the field $\mathbb{F}_{q}$ there are no other orbits. The main result now is completely proved.
Similarly, other root systems $\Phi$ can be considered. Let us fix a linear lexicographic order $\succ$ on the positive roots of $\Phi^{+}$. Let $\lambda\in\mathfrak{n}^{*}$, $\gamma\in\Phi^{+}$. Consider the following matrices.
@@
Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: any $\gamma\in\textup{supp}(\lambda)$ satisfies condition (1). We can formulate a hypothesis.
Conjecture 3.8. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.

### v4

@@ 3. Classification of orbits
The sum of the coefficients is exactly 883, which is the number of supports in the list plus three supports with restrictions.
Since each form in $S$ determines its own orbit, the set $S$ accounts for $|S|$ orbits. However, in the work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$. Hence, over $\mathbb{F}_{q}$ there are no other orbits. This completes the proof of the main result.
The same approach can be applied to other root systems $\Phi$. Fix a total lexicographic order $\succ$ on the positive roots $\Phi^{+}$. For $\lambda\in\mathfrak{n}^{*}$ and $\gamma\in\Phi^{+}$, define the matrices
@@
Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\textup{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
Conjecture 3.8. For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.